From mar@roe.ac.uk Tue Jun 25 15:22:52 2013 Date: Fri, 21 Jun 2013 14:12:54 +0100 From: Mike Read To: VDFS Science Archive Team -- Clive Davenhall , Eckhard Sutorius , Mark Holliman , Nicholas Cross , Nigel Hambly , Rob Blake , Ross Collins , Stelios Voutsinas Cc: CCs for VDFS -- Alastair Edge , Andrew Lawrence , Bob Mann , Dante Minniti , James Dunlop , Jim Emerson , Keith Noddle , Lorenzo Rimoldini , Maria-Rosa Cioni , Matt Jarvis , Mike Irwin , Nigel Metcalfe , Norman Gray , Phil Lucas , Richard McMahon at IOA , Simon Dye , Stephen Warren , Tom Kerr , Tom Shanks Subject: WFAU VDFS science archives meeting minutes, 14/6/13 Present: RPB, RGM, NJC, MSH, MAR Apologies: JPE, AL, STV, NG, KTN, ACD, ETWS, NCH, RSC DoNM: 10am Friday June 28th 2013 in the Villas meeting room NCH in the chair (room is booked). Actions discharged this time: ----------------------------- ACTION: NJC to check the standard crossmatching of non-survey PI programmes for inclusion of WISE (all) and SDSS-DR9 (WFCAM) Done. ACTION: NJC to update VISTA auto-QC to deprecate C/D/R flagged frames Done. Actions partly discharged but continuing: ----------------------------------------- ACTION: MAR to have a go at releasing other VISTA PI DBs (e.g. for JPE) Continues; curation code exiting with errors. Actions carried forward from 31/05/2013 meeting ----------------------------------------------- ACTION: ETWS, with RPB, to develop an automated helper script to keep release DB files links updated. Continues; this one still has low priority. ACTION: NCH to enquire with MAST as to the status of GALEX GR6 & 7 and see if an exchange of SQL DBs is warranted. Continues. Specific points and new actions: -------------------------------- Project management: WFAU quarterly to be held Friday 21st June. WFCAM, VISTA and VST updates: Nothing to report. Comments and issues arising from CASU minutes: No new minutes this time. Networking: Slurping over of VISTA v1.3 continues, currently transferring P88. WSA/VSA/OSA/GES Operations: Mainly VISTA v1.3 curation. Notifications of GES community contacts have been coming in, access has been granted and they have been setting up accounts. Hardware and Systems: The new NAS box, Shebaka, has arrived. The C: drive on Ramses5 was showing as nearly full, however it wasn't obvious what files were causing this as when everything is added up there should be plenty of space. Survey Data Release: The much revised VMC ESO release had been re-generated and uploaded to ESO. The ATLAS release with new UI was now scheduled for mid-July. NJC had tracked down the problem with inconsistent VMC photometry across releases. The summary being that VMCv20121128 has erroneous zero-point offsets of the order of 0.02 mags and some of these have been propagated into VMCv20130304 (re-used deep stacks). NJC has emailed details of the issue to the VMC team. ACTION: NJC to add a note to VSA knownIssues detailing VMC photometry problem. The VMC team would like to see a new release that fixes the issue but also to include P90 data. Our intention had been to use only v1.3 data for P90 and any other VISTA releases going forward. However given the timetable for having all the v1.3 data transferred and ingested it was agreed to attempt a fixed and updated v1.2/1.3 VMC P90 release. Although this would entail some duplication of curation and data on the two VSA load servers it was thought that the system and software could cope. The astrometry problem in GPS DR8 had been tracked down to inconsistent WCS keywords (PV2_3 and PROJP3) in the re-processed data. This affects around 3% of sources in the gpsSource. Mike Irwin had identified all the effected files (05A-06A). Details had been emailed to Phil Lucas. A note would be added to the knownIssues webpage and the files would be fixed or deprecated in the next release. ACTION: MAR to add a note to WSA knownIssues detailing GPS astrometry problem in DR8. Software: NJC reported work on the VMC photometry fix and ESO grade cut for the nonSurvey code. Non-survey Data Release: Eight nonSurvey release DBs had been created. Astrogrid deployment, VO & Data Analysis services: MSH noted he'd created UKIDSS DR5 MOC coverage maps at a user's request and had fixed some of the metadocs for the SDSS TAP services. Miscellaneous: Stelios and Dave would at ROE the following week (Tues-Fri). -- Scanned by iCritical.